OpenStudy (anonymous):

what is the domain of 5 - (x+20)^2

OpenStudy (cgreenwade2000):

Because x can equal any real number, the answer is all real numbers.

OpenStudy (anonymous):

thank you find the x and y intercepts

OpenStudy (cgreenwade2000):

>_> Write the whole problem. Right now there is no y.

OpenStudy (anonymous):

f(x) = 5 - (x + 20)^2

OpenStudy (cgreenwade2000):

f(x) is the same as writing y. So, y= 5- (x+20)^2 Now, what is the value of y on the x-axis?

OpenStudy (anonymous):

idk. it just says that y=f(x) in the xy plane

OpenStudy (cgreenwade2000):

1. Make a graph. 2. Look at the x-axis 3. What is the value of y at any point on the x-axis?

OpenStudy (anonymous):

well, I know it's a parabola and that the vertex is -20, 5

OpenStudy (cgreenwade2000):

>_> On the x-axis, y = 0 On the y-axis, x=0 So to find the x intercept, we set y = 0 To find the y intercept, we set x = 0 Tell me what you get for each after you have solved for both intercepts.

OpenStudy (anonymous):

oh! y = -395

OpenStudy (cgreenwade2000):

That's the y-intercept. Find the x-intercept.

OpenStudy (anonymous):

that's what I meant. finding the x intercept now

OpenStudy (cgreenwade2000):

Hint: x-intercept(s)

OpenStudy (anonymous):

yes, because it's a parabola, right?

OpenStudy (anonymous):

I'm having difficulty with this. I know part of it is -20

OpenStudy (anonymous):

what do I do with the 5?

OpenStudy (anonymous):

I have (x+20)^2=5